# Complex-valued in-medium potential between heavy impurities in ultracold atoms

### v1

@@ I Introduction
Generally, however, such real-valued interaction potentials cannot describe all of the relevant medium effects. Polaron physics is inherently an open quantum system problem, possessing non-Hermitian nature due to its environmental medium effect. Recently, the open quantum system problem has been believed to be relevant in broad fields of science, ranging from quantum computations [40, 41], topological quantum matters [42, 43], to photosynthetic biological materials [44, 45]: coupling to an environment not only can lead to decoherence and dissipation of the system but also can bring about entirely new emergent phenomena. To elucidate the open-system nature of polaron physics, it is essential to reconsider the medium-induced interaction, which is a fundamental building block of the many-body physics of polarons; due to the non-Hermitian character, it is naturally expected that the induced potential becomes complex-valued. The complex-valued potential between the heavy quarks [46, 47, 48, 49] have also been believed to be important in clarifying heavy-quark dynamics in quark-gluon plasmas—a hot QCD matter created during heavy-ion collision experiments.
@@
To address this question, we study in this paper the complex-valued in-medium potential between the two heavy polarons in cold atomic systems, especially focusing on its imaginary part. Following the definition developed in the context of hot QCD plasmas [46, 47, 48, 49], we lay out a solid theoretical basis of the complex-valued in-medium potential suited for the study of cold atomic systems. We also present a practical formula for the complex-valued potential when the impurity-medium coupling is weak, which was first derived in [50] in the context of cold atomic gases using the influence functional formalism. We then evaluate the imaginary part of the in-medium potential for polarons immersed in two representative examples of the finite-temperature quantum medium—a Fermi gas and a superfluid gas. Analogous to the real part, we identify that it shows an oscillating-decaying behavior for the Fermi polarons at low temperatures while it exhibits a monotonic decay for the polarons in the superfluid. Notably, we find the long-range asymptotic behavior shows the same power-law decay given by $V_{\mathrm{Im}}(r)\propto 1/r^{2}$ in both cases. We stress that this is not a coincidence but rather a universal result: we show that the $1/r^{2}$ imaginary potential universally appears for any physical system if the collision between the polaron and an elementary excitation in the medium is almost elastic. We also discuss three experimental manifestations of the imaginary part of the complex-valued potential; that is, a direct detection through radio-frequency interferometry, and relatively indirect methods via spectral width of bipolaron, and relaxation of density fluctuation induced by the single impurity.
The paper is organized as follows. In Sec. II, we provide a general definition of the in-medium potential and formulae to evaluate the real and imaginary parts applicable when the impurity and medium is weakly coupled through a $s$-wave channel. In Sec. III, we evaluate the imaginary part of the in-medium potential acting on two polarons immersed in a non-interacting Fermi gas and a superfluid gas. In Sec. IV, we discuss experimental protocols to observe the imaginary potential. Sec. V is devoted to the conclusion and future outlook. In Appendices A and B, we provide detailed derivations of some results in the main text, and the relation between our work and the Langevin equation, respectively. Throughout this paper, a natural unit $\hbar=k_{B}=1$ is used.
@@ II In-medium potential
In this section, we lay out the basis of the in-medium potential between the two impurities immersed in a finite-temperature medium relying on the real-time Green’s function, following Ref. [46]. After presenting a general definition of the in-medium potential in Sec. II.1, we derive a useful formula for the weakly-coupled impurity-medium case in Sec. II.2, which will be used in the subsequent sections.
II.1 Definition
Let us consider two impurities with their masses $M$ immersed in a finite-temperature Fermi or Bose medium. To define the medium-induced potential acting on the impurities, we rely on an impurity two-body correlation function, which we can regard as a generalization of the wave function for the relative coordinates of the two impurities in the medium. Using the impurity creation operator $\hat{\Phi}^{\dagger}(\bm{x})$ at position $\bm{x}$, we introduce the following real-time correlation function:
@@ II.1 Definition
×$\times\hat{\Phi}^{{\dagger}}(\bm{x}_{1})\hat{\Phi}^{{\dagger}}(\bm{x}_{2})\big{\rangle},$ (1)
[…]
@@
where we introduced the energy of two impurities in the medium $\bar{V}(\bm{r},\bm{\nabla}_{r})$ subtracted by the kinetic energy with an in-medium effective impurity mass $M_{*}$. We note that $\bar{V}(\bm{r},\bm{\nabla}_{r})$ generally contains spatial derivatives due to impurity motions. In the limit where the impurity has a sufficiently heavy mass, we can treat the impurity as a test particle almost fixed at a certain position, and thus drop the derivative dependence of the kinetic term and $\bar{V}(\bm{r},\bm{\nabla}_{r})$. As a result, in this limit, the time-derivative and the potential only survive in Eq. (2). We then define the in-medium energy for two infinitely heavy impurities as $\bar{V}(\bm{r})$ from the following ratio of the impurity correlation function (1):
@@
As we will explicitly see later, the presence of the finite-temperature medium makes the induced potential $V(\bm{r})$ complex-valued as
@@ II.2 Weak impurity-medium coupling case
where $g$ is the impurity-medium coupling constant. One can relate this coupling constant to the $s$-wave scattering length $a_{\mathrm{IM}}$ between the impurity and medium particles as $g=2\pi a_{\mathrm{IM}}\left(\frac{1}{m}+\frac{1}{M}\right)$ with masses $m$ and $M$ for the medium and impurity particles [51, 52]. We also introduced the particle number density operator for the medium $\hat{n}(\bm{x})$.
For a system with weak impurity-medium coupling, the interaction picture with the choice of the unperturbed Hamiltonian
with $\mathrm{T}$ denoting the time-ordered product. Here, we introduced $\hat{\mathcal{O}}(t)\equiv\mathrm{e}^{\mathrm{i}t\hat{H}_{0}}\hat{\mathcal{O}}\mathrm{e}^{-\mathrm{i}t\hat{H}_{0}}$ for an arbitrary operator $\hat{\mathcal{O}}$, and the number density fluctuation operator around its thermal average as $\delta\hat{n}(\bm{x})=\hat{n}(\bm{x})-\langle\hat{n}(\bm{x})\rangle$. Note that the time dependence of the number density fluctuation operator originates solely from the medium Hamiltonian $\hat{H}_{\mathrm{med}}$ as $\delta\hat{n}(\bm{x},t)=\mathrm{e}^{\mathrm{i}t\hat{H}_{\mathrm{med}}}\delta\hat{n}(\bm{x})\mathrm{e}^{-\mathrm{i}t\hat{H}_{\mathrm{med}}}$ because only $\hat{H}_{\mathrm{med}}$ in $\hat{H}_{0}$ contains the medium field.
We take the heavy-impurity limit as the first step to evaluate $\Psi(\bm{r},t)$. In this limit, the large impurity mass $M_{*}^{-1}\simeq M^{-1}$ suppresses the impurity kinetic terms. Combined with the fact that the self-interactions between the impurities are negligible for a few impurities, we can ignore $\hat{H}_{\mathrm{imp}}$ as
$\hat{H}_{0}\simeq\hat{H}_{\mathrm{med}}+g\int\mathrm{d}^{3}x\,\hat{\Phi}^{\dagger}(\bm{x})\hat{\Phi}(\bm{x})\langle\hat{n}(\bm{x})\rangle.$ (10)
The vital point here is that Eq. (10) commutes with the $\hat{\Phi}^{\dagger}(\bm{x})\hat{\Phi}(\bm{x})$. This allows us to drop the time-dependence of the impurity field operator in Eq. (9) as
$\hat{U}(t,0)\simeq\mathrm{e}^{-\mathrm{i}tH_{0}}\mathrm{T}\mathrm{e}^{-\mathrm{i}g\int^{t}_{0}\mathrm{d}t^{\prime}\mathrm{d}^{3}x\,\hat{\Phi}^{\dagger}(\bm{x})\hat{\Phi}(\bm{x})\delta\hat{n}(\bm{x},t^{\prime})},$ (11)
meaning that the impurity is immovable. The above procedure clarifies a precise definition of the heavy-impurity approximation, in which we treat the impurity as a test particle fixed at a certain position.
Substituting Eq. (11) into Eq. (1), we find $\Psi(\bm{r},t)$ to be
To proceed further, we employ the weak-coupling expansion with respect to the impurity-medium coupling $g$. As a ladder resummation, we expand the exponential function in the average and approximate the correlation functions by factorizing them into two point functions of $\delta\hat{n}$. Then we obtain
where we used the translational and rotational invariance of the medium. Matching this expression with the aforementioned long-time behavior $\Psi(\bm{r},t)\simeq\mathrm{e}^{-\mathrm{i}\bar{V}(\bm{r})t}\Psi(\bm{r},0)$, we find $\bar{V}(\bm{r})$ at $O(g^{2})$ as
In this result, the $\bm{r}$-independent contribution survives even when two impurities are infinitely separated. Assuming a cluster decomposition property, we identify this part as twice the energy for a single impurity immersed in the medium. According to Eq. (4), subtracting this one-body energy, we arrive at the expression of the in-medium interaction potential $V(\bm{r})$ for a weak impurity-medium coupling case as
$V(\bm{r})$ $\simeq\lim_{t\to\infty}\frac{g^{2}}{\mathrm{i}t}\int_{0}^{t}\mathrm{d}t_{1}\int_{0}^{t}\mathrm{d}t_{2}\,\langle\mathrm{T}\delta\hat{n}(\bm{r},t_{1})\delta\hat{n}(\bm{0},t_{2})\rangle,$
$=\lim_{t\to\infty}\frac{2g^{2}}{\mathrm{i}t}\int_{0}^{t}\mathrm{d}s\,(t-s)\langle\delta\hat{n}(\bm{r},s)\delta\hat{n}(\bm{0},0)\rangle$
where we dropped $O(g^{3})$ terms. In the second line, noting that the correlation function only depends on the difference of two-time arguments, we perform the one time integral. To obtain the third line, we further neglected the term proportional to $s$ since the correlation function $\langle\delta\hat{n}(\bm{r},s)\delta\hat{n}(\bm{0},0)\rangle$ at fixed $\bm{r}$ is expected to decay in a finite duration of time.
By separating the real and imaginary parts of Eq. (15), we can express them with the retarded Green’s function, defined by the number density fluctuation in the position-frequency representation as
$G^{R}(\bm{r},\omega)$ $\equiv\mathrm{i}\int_{-\infty}^{\infty}\mathrm{d}t\,\mathrm{e}^{\mathrm{i}\omega t}\theta(t)\langle[\delta\hat{n}(\bm{r},t),\delta\hat{n}(\bm{0},0)]\rangle,$ (16)
where $\theta(t)$ is the Heaviside step function. Recalling $\langle\hat{A}\hat{B}\rangle^{*}=\langle\hat{B}^{\dagger}\hat{A}^{\dagger}\rangle$, we find the real and imaginary parts of Eq. (15) as
@@
where we used the fluctuation-dissipation relation
$V_{\mathrm{Im}}(\bm{r})$ $=-\frac{g^{2}}{2}\int_{-\infty}^{\infty}\mathrm{d}s\,\Big{\langle}\delta\hat{n}(\bm{r},s)\delta\hat{n}(\bm{0},0)+\delta\hat{n}(\bm{0},0)\delta\hat{n}(\bm{r},s)\Big{\rangle}$
$=-g^{2}\lim_{\omega\to 0}\frac{2T}{\omega}\mathop{\mathrm{Im}}G^{R}(\bm{r},\omega).$ (19)
Equations (17) and (18) reproduce the formula given in [50], whose main subject was classical Langevin equation. We provide a brief explanation in Appendix B on how the complex potential and the classical Langevin equation are related. Recalling Eq. (14), we also find the $\bm{r}$-independent one-body energy $E^{(1)}=E_{\mathop{\mathrm{Re}}}^{(1)}+\mathrm{i}E_{\mathop{\mathrm{Im}}}^{(1)}$ as
@@
In summary, we obtain practically useful formulae (17) and (18) in terms of the medium retarded Green’s function (16). In the following sections, we derive the in-medium potential by computing the retarded Green’s function with the use of the imaginary-time formalism.
Some remarks are in order. First, in the weak impurity-medium coupling regime, the in-medium potential $V(\bm{r})$ is determined by the dynamical properties of the medium without the impurities [recall that the time-evolution of $\delta\hat{n}(\bm{x},t)$ is controlled by $\hat{H}_{\mathrm{med}}$]. Making use of this viewpoint, we will give a proposal to experimentally investigate an imprint of the in-medium potential on medium dynamics with a single impurity (see Sec. IV.3). Second, the obtained formula for the in-medium potential (15) provides a natural generalization of the well-known formula at $T=0$. In fact, using the time-ordered product, one finds Eq. (15) to be
$V(\bm{r})=-\mathrm{i}g^{2}\int_{-\infty}^{\infty}\mathrm{d}t\,\langle\mathrm{T}\delta\hat{n}(\bm{r},t)\delta\hat{n}(\bm{0},0)\rangle.$ (22)
In the zero temperature limit, the thermal average reduces to the ground state average, yielding a familiar result at weak coupling [53]. Finally, Eq. (18) is proportional to the temperature $T$. As a result, the imaginary part representing the dissipative effect is peculiar to finite-temperature media.
@@ III.1 Polarons in Fermi gas
where $m$ and $\mu$ denote the mass and chemical potential of the medium fermions. In the present setup, we can exactly compute Eqs. (17) and (18) thanks to the non-interacting property of the single-component fermion.
To compute the retarded Green’s function (16), we rely on the imaginary-time formalism and perform the analytic continuation. For that purpose, let us introduce the Matsubara Green’s function for the medium fermion as
[…]
where we introduced the fermionic Matsubara frequency $\omega_{n}^{F}=2\pi(n+\frac{1}{2})T$ with a medium temperature $T$ and the energy measured from the Fermi sea $\xi_{\bm{k}}=\frac{\bm{k}^{2}}{2m}-\mu$.
The Matsubara Green’s function for the number density operator $\hat{n}=\hat{\psi}_{f}^{\dagger}\hat{\psi}_{f}$ corresponding to Eq. (16) is given by the diagram shown in Fig. 1 with the impurity legs amputated and reads
$G(\bm{k},\mathrm{i}\omega_{n}^{B})$ $=T\sum_{l}\int_{\bm{q}}S(\mathrm{i}\omega_{n}^{B}+\mathrm{i}\omega_{l}^{F},\bm{k}+\bm{q})S(\mathrm{i}\omega_{l}^{F},\bm{q})$
$=-\int_{\bm{q}}\frac{n_{F}(\xi_{\bm{k}+\bm{q}})-n_{F}(\xi_{\bm{q}})}{\mathrm{i}\omega_{n}^{B}-\xi_{\bm{k}+\bm{q}}+\xi_{\bm{q}}},$ (25)
where we defined the Matsubara frequency for bosons $\omega^{B}_{n}=2\pi nT$, the Fermi distribution function $n_{F}(E)\equiv 1/(\mathrm{e}^{\beta E}+1)$, and a shorthand notations for the integral as $\int_{\bm{k}}\equiv\int\mathrm{d}^{3}k/(2\pi)^{3}$.
@@
We then perform the analytic continuation and the Fourier transformation to obtain the retarded Green’s function in real space as
@@
Using Eqs. (17) and (18) together with the formula $\frac{1}{x+\mathrm{i}0}=\mathrm{P}\frac{1}{x}-\mathrm{i}\pi\delta(x)$, we find
@@
×$\hskip 40.0pt\times\frac{1}{T}n_{F}(\xi_{\bm{k}})\big{[}1-n_{F}(\xi_{\bm{k}})\big{]}.$ (28)
Note that $V_{\mathop{\mathrm{Re}}}(\bm{r})$ gives a finite-temperature real potential, which reduces to the RKKY interaction at zero temperature. Our calculation reproduces the complex-valued potential in the same setup obtained by influence functional formalism [50].
Performing the remaining integrals for $V_{\mathop{\mathrm{Im}}}(\bm{r})$, we obtain the following result:
@@
where we neglected the higher-order correction with respect to $k_{F}r$ or $(k_{F}r)^{-1}$. We note that the leading short-distance value in Eq. (31) gives the imaginary part of the one-body energy [recall Eq. (21)]. On the other hand, in the low-temperature limit $T/T_{F}\ll 1$, the Fermi distribution reduces to the Heaviside step function, and we obtain
@@
with $f(y,0)=(1-\cos y)/y^{2}$. Similarly to the RKKY interaction, the imaginary part of the potential shows an oscillatory behavior in the low-temperature limit, whose period is controlled by the Fermi momentum $k_{F}$. As we mentioned, $V_{\mathop{\mathrm{Im}}}(\bm{r})$ vanishes at $T=0$ because the magnitude is proportional to the temperature $T$. Moreover, the exponent of the power-law decay is $V_{\mathop{\mathrm{Im}}}(\bm{r})\propto r^{-2}$, which contrasts a difference with that for the RKKY interaction, $V_{\mathop{\mathrm{Re}}}(\bm{r})\propto r^{-4}$.
Figure 2 demonstrate $V_{\mathop{\mathrm{Im}}}(\bm{r})$ for several temperatures normalized by either $T_{F}$ or $T$. The dashed curve in Fig. 2 (a)—corresponding to the zero-temperature limit—demonstrates the oscillatory decaying around $T=0$ against the inter-polaron separation $r$. This oscillatory behavior is suppressed when we increase temperature. On the other hand, the magnitude of $V_{\mathop{\mathrm{Im}}}(\bm{r})$ is linearly increasing with respect to $T$. This is because the imaginary part corresponds to the dissipative effect associated with the finite-temperature medium. Figure 2 (b) also demonstrates the $r^{-2}$ decay at long distances specified in Eq. (32).
Refer to caption (a)
Refer to caption (b)
Figure 2: (a) The imaginary part of the induced potential for Fermi polarons with $T/T_{F}=0.1,0.5$, and $1.0$ (blue, green, and red curves), normalized by $T$. The dashed curve shows the zero-temperature limit of the imaginary potential. (b) The imaginary part of the induced potential at long distances for Fermi polarons with $T/T_{F}=0.1,0.5$, and $1.0$ (blue, green, and red curves), normalized by $T_{F}$. The dashed curve shows the power-law decay $r^{-2}$, which matches those of the induced potential.

### v2

@@ I Introduction
Generally, however, such real-valued interaction potentials cannot describe all of the relevant medium effects. Polaron physics is inherently an open quantum system problem, possessing non-Hermitian nature due to its environmental medium effect. Recently, the open quantum system problem has been believed to be relevant in broad fields of science, ranging from quantum computations [64, 7], topological quantum matters [6, 30], to photosynthetic biological materials [33, 65]: coupling to an environment not only can lead to decoherence and dissipation of the system but also can bring about entirely new emergent phenomena. To elucidate the open-system nature of polaron physics, it is essential to reconsider the medium-induced interaction, which is a fundamental building block of the many-body physics of polarons; due to the non-Hermitian character, it is naturally expected that the induced potential becomes complex-valued. The complex-valued potential between the heavy quarks [43, 5, 10, 67] have also been believed to be important in clarifying quark dynamics of quark-gluon plasmas—a hot QCD matter created during heavy-ion collision experiments.
@@
To address this question, we study in this paper the complex-valued in-medium potential between the two heavy polarons in cold atomic systems, especially focusing on its imaginary part. Following the definition developed in the context of hot QCD plasmas [43, 5, 10, 67], we first lay out a solid theoretical basis of the complex-valued in-medium potential suited for the study of cold atomic systems, which is applicable even in the case with strong impurity-medium coupling. We also present a practical formula for the complex-valued potential when the impurity-medium coupling is weak, which was first derived in [73] in the context of cold atomic gases using the influence functional formalism. We then evaluate the imaginary part of the in-medium potential for polarons immersed in two representative examples of the finite-temperature quantum medium—a Fermi gas and a superfluid gas. Analogous to the real part, we identify that it reproduced an oscillating-decaying behavior for the Fermi polarons at low temperatures [73] while it exhibits a monotonic decay for the polarons in the superfluid. Notably, we find the long-range asymptotic behavior shows the same power-law decay given by $V_{\mathrm{Im}}(r)\propto 1/r^{2}$ in both cases. We stress that this is not a coincidence but rather a universal result: we show that the $1/r^{2}$ imaginary potential universally appears for any physical system if the collision between the polaron and an elementary excitation in the medium is almost elastic. We also discuss three experimental manifestations of the imaginary part of the complex-valued potential; that is, a direct detection through radio-frequency interferometry, and relatively indirect methods via spectral width of bipolaron, and relaxation of density fluctuation induced by the single impurity.
The paper is organized as follows. In Sec. II, we provide a general definition of the in-medium potential and formulae to evaluate the real and imaginary parts applicable when the impurity and medium is weakly coupled through a $s$-wave channel. In Sec. III, we evaluate the imaginary part of the in-medium potential acting on two polarons immersed in a non-interacting Fermi gas and a superfluid gas and discuss the physical origin of its universal $r^{-2}$ behavior. In Sec. IV, we discuss experimental protocols to observe the imaginary potential. Sec. V is devoted to the conclusion and future outlook. In Appendices A, B, and C, we provide detailed derivations of some results in the main text, and the relation between our work and the Langevin equation. Throughout this paper, a natural unit $\hbar=k_{B}=1$ is used.
@@ II In-medium potential
In this section, we lay out the basis of the in-medium potential between the two impurities immersed in a finite-temperature medium relying on the real-time Green’s function, following Ref. [43]. After presenting a general definition of the in-medium potential, we give a formula for the weakly-coupled impurity-medium case reproducing the one derived in Ref. [73], which will be used in the subsequent sections.
Consider two impurities with their masses $M$ immersed in a finite-temperature Fermi or Bose medium. To define the medium-induced potential acting on the impurities, we rely on an impurity two-body correlation function, which we can regard as a generalization of the wave function for the relative coordinates of the two impurities in the medium. Using the impurity creation operator $\hat{\Phi}^{\dagger}(\bm{x})$ at position $\bm{x}$, we introduce the following real-time correlation function:
@@
$\times\hat{\Phi}^{{\dagger}}(\bm{x}_{1})\hat{\Phi}^{{\dagger}}(\bm{x}_{2})\big\rangle,$ (1)
where $\langle\cdots\rangle$ denotes the average over the density operator $\hat{\rho}_{0}=\hat{\rho}_{\mathrm{eq}}^{\mathrm{med}}\otimes|0_{\mathrm{imp}}\rangle\langle 0_{\mathrm{imp}}|$ with an equilibrium density operator for the medium $\hat{\rho}_{\mathrm{eq}}^{\mathrm{med}}$ and the zero-particle state for the impurity $|0_{\mathrm{imp}}\rangle$, and $\hat{U}(t,0)$ is a time-evolution operator of the total system from 0 to $t$. We let the initial condition, or $\bm{x}_{1}$ and $\bm{x}_{2}$ dependence, of the wave function implicit. One can then confirm that $\Psi(\bm{r},t)$ reduces to the usual wave function for the relative motion of two impurities separated by a distance $\bm{r}$ in the absence of the surrounding thermal medium.
@@
where we introduced the energy of two impurities in the medium $\bar{V}(\bm{r},\bm{\nabla}_{r})$ subtracted by the kinetic energy with an in-medium effective impurity mass $M_{*}$. We note that $\bar{V}(\bm{r},\bm{\nabla}_{r})$ generally contains spatial derivatives due to impurity motions. In the limit where the impurity is sufficiently heavy, we can treat the impurity as a test particle almost fixed at a certain position, and thus drop the derivative dependence of the kinetic term and $\bar{V}(\bm{r},\bm{\nabla}_{r})$. As a result, the time-derivative and the potential only survive in Eq. (2). We can then define the in-medium energy for two infinitely heavy impurities as follows:
@@
As shown later, the presence of the finite-temperature medium makes the induced potential complex-valued as
@@
where $g$ is the impurity-medium coupling constant. One can relate $g$ to the $s$-wave scattering length $a_{\mathrm{IM}}$ between the impurity and medium particles as $g=2\pi a_{\mathrm{IM}}\left(\frac{1}{m}+\frac{1}{M}\right)$ with masses $m$ and $M$ for the medium and impurity particles [9, 63]. We also introduced the particle number density operator for the medium $\hat{n}(\bm{x})$.
For a weak impurity-medium coupling $g$, the real-time correlation function $\Psi(\bm{r},t)$ can be computed by the ladder approximation. Using $\Psi(\bm{r},t)\simeq\mathrm{e}^{-\mathrm{i}\bar{V}(\bm{r})t}\Psi(\bm{r},0)$ and Eq. (4), we obtain the complex-valued potential as follows (see Appendix A for details):
@@
where we introduced the retarded Green’s function of the number density
$G^{R}(\bm{r},\omega)$ $\equiv\mathrm{i}\int_{-\infty}^{\infty}\mathrm{d}t\,\mathrm{e}^{\mathrm{i}\omega t}\theta(t)\langle[\hat{n}(\bm{r},t),\hat{n}(\bm{0},0)]\rangle,$ (10)
with $\hat{n}(\bm{r},t)\equiv\mathrm{e}^{\mathrm{i}\hat{H}_{\textrm{med}}t}\hat{n}(\bm{r})\mathrm{e}^{-\mathrm{i}\hat{H}_{\textrm{med}}t}$ and the Heaviside step function $\theta(t)$. We also find the $\bm{r}$-independent one-body energy $E^{(1)}=E_{\mathop{\mathrm{Re}}}^{(1)}+\mathrm{i}E_{\mathop{\mathrm{Im}}}^{(1)}$ as
@@
We note that Eqs. (8) and (9) reproduce the formula given in [73], whose main subject was classical Langevin equation. We provide a brief explanation in Appendix B on how the complex potential and the classical Langevin equation are related.
Some remarks are in order. First, in the weak impurity-medium coupling regime, the in-medium potential $V(\bm{r})$ is determined by the dynamical properties of the medium without the impurities [recall that the time-evolution of $\hat{n}(\bm{x},t)$ is controlled by $\hat{H}_{\mathrm{med}}$]. We will later use this to propose a way to experimentally investigate an imprint of the in-medium potential on medium dynamics with a single impurity (see Sec. IV.3). Second, Eq. (9) is proportional to the temperature $T$. As a result, the imaginary part representing the dissipative effect is peculiar to finite-temperature media. Finally, as we show in the Appendix A, the derived formula provides a natural generalization of the well-known formula at $T=0$.
@@ III.1 Polarons in Fermi gas
where $m$ and $\mu$ denote the mass and chemical potential of the medium fermions.
@@
For non-interacting fermions, we can exactly compute the Green’s function from the diagram in Fig. 1 as
$G(\bm{k},\mathrm{i}\omega_{n}^{B})$ $=-\int_{\bm{q}}\frac{n_{F}(\xi_{\bm{k}+\bm{q}})-n_{F}(\xi_{\bm{q}})}{\mathrm{i}\omega_{n}^{B}-\xi_{\bm{k}+\bm{q}}+\xi_{\bm{q}}},$ (14)
with the bosonic Matsubara frequency $\omega^{B}_{n}=2\pi nT$ at temperature $T$, Fermi distribution $n_{F}(E)\equiv 1/(\mathrm{e}^{E/T}+1)$, the energy measured from the Fermi sea $\xi_{\bm{k}}=\frac{\bm{k}^{2}}{2m}-\mu$, and a shorthand notations for the integral $\int_{\bm{k}}\equiv\int\mathrm{d}^{3}k/(2\pi)^{3}$. Performing the analytic continuation and the Fourier transformation, the retarded Green’s function is obtained as
@@
Using Eqs. (8) and (9), we find
@@
$\hskip 40.0pt\times\frac{1}{T}n_{F}(\xi_{\bm{k}})\big[1-n_{F}(\xi_{\bm{k}})\big].$ (17)
$V_{\mathop{\mathrm{Re}}}(\bm{r})$ gives a finite-temperature real potential, which reduces to the RKKY interaction at $T=0$. Our calculation reproduces the complex-valued potential in the same setup obtained by influence functional formalism [73].
Performing the integrals for $V_{\mathop{\mathrm{Im}}}(\bm{r})$, we obtain
@@
where we neglected the higher-order correction with respect to $k_{F}r$ or $(k_{F}r)^{-1}$. We note that the leading short-distance value in Eq. (20) gives the imaginary part of the one-body energy [recall Eq. (12)]. On the other hand, in the low-temperature limit, we obtain
@@
with $f(y,0)=(1-\cos y)/y^{2}$. Similarly to the RKKY interaction, the imaginary part of the potential shows an oscillatory behavior in the low-temperature limit, whose period is controlled by the Fermi momentum $k_{F}$. $V_{\mathop{\mathrm{Im}}}(\bm{r})$ vanishes at $T=0$ because it is proportional to $T$. Moreover, the exponent of the power-law decay is $V_{\mathop{\mathrm{Im}}}(\bm{r})\propto r^{-2}$, which contrasts a difference with that for the RKKY interaction, $V_{\mathop{\mathrm{Re}}}(\bm{r})\propto r^{-4}$.
Figure 2 shows $V_{\mathop{\mathrm{Im}}}(\bm{r})$ for several temperatures normalized by either $T_{F}$ or $T$. The dashed curve in Fig. 2 (a)—corresponding to the zero-temperature limit—demonstrates the oscillatory decay around $T=0$ against the inter-polaron separation $r$. This oscillation is suppressed at higher temperature. On the other hand, the magnitude of $V_{\mathop{\mathrm{Im}}}(\bm{r})$ increases linearly as $T$. This is because the imaginary part corresponds to the dissipative effect associated with the finite-temperature medium. Figure 2 (b) also demonstrates the $r^{-2}$ decay at long distances specified in Eq. (21).
Figure 2: (a) The imaginary part of the induced potential for Fermi polarons with $T/T_{F}=0.1,0.5$, and $1.0$ (solid blue, dashed-dotted green, and dashed red curves), normalized by $T$. The dotted black curve shows the zero-temperature limit of the imaginary potential. (b) The imaginary part of the induced potential at long distances for Fermi polarons with $T/T_{F}=0.1,0.5$, and $1.0$ (solid blue, dashed-dotted green, and dashed red curves), normalized by $T_{F}$. The dotted black curve shows the power-law decay $r^{-2}$, which matches those of the induced potential.